Solve (D^2-D)z = cos(3x-y)
1
Expert's answer
2020-11-02T20:03:05-0500
"D(D-1)z=\cos(3x-y)"

"m_1=0, m_2=0, c_1=0, c_2=1"

Complementary function is "f_1(y)+e^xf_2(y)."

Let "z=A\sin(3x-y)+B\cos(3x-y)"



"-9A\sin(3x-y)-9B\cos(3x-y)-"

"-(3A\cos(3x-y)-3B\sin(3x-y))=\cos(3x-y)"

"\begin{matrix}
 -9A+3B=0 \\
 -9B-3A=1
\end{matrix}"

"\begin{matrix}
 B=3A \\
 -27A-3A=1
\end{matrix}"

"A=-{1 \over 30}"

"B=-{1 \over 10}"

"P.I.=-\dfrac{1}{30}\sin(3x-y)-\dfrac{1}{10}\cos(3x-y)"

Hence the complete solution is 

"z=f_1(y)+e^xf_2(y)-\dfrac{1}{30}\sin(3x-y)-\dfrac{1}{10}\cos(3x-y)"